Isaac Diez. Isaac Diez Peris lives in Rio de Janeiro. While attending school in Spain he meets Juan Carlos Cordero from Guatemala. Over the summer holiday Isaac decides to visit Juan Carlos in Guatemala City for a couple of weeks. Isaacs parents give him some spending money, R$4,900. Isaac wants to exchange it for Guatemalan quetzals (GTQ). He collects the following rates Spot rate on the GTQ/E exchange rate is GTQ10.6482- 1.00 Spot rate on the /RS exchange rate is 0.3675 R$1.00 a. What is the Brazilian real/Guatemalan quetzal cross rate? b. How many quetzals will Isaac get for his reais?
Answer #1

a) Spot rate GTQ/Euro = 10.6842/1
Spot Rate Euro/Real = 0.3675/1
So 1 Euro = 1/0.3675 Real = 2.7211
So Brazil Real/Guatemalan Quetzal cross rate = 2.7211/10.6482 = 0.2555 Brazil Real/Guatemalan Quetzal

b) Amount of Quetzal for 4900 real = 4900* 1/0.2555 = 19178.08


Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well